# PGT-I: Scaling Spatiotemporal GNNs with Memory-Efficient Distributed Training

### v2

@@ 2. Background
This work focuses on the distributed training of ST-GNNs. Accordingly, section 2.1 provides an introduction to spatiotemporal data, section 2.2 introduces ST-GNNs and DCRNN, and section 2.3 details ST-GNN data preprocessing. To provide clarity, we define the terminology relating to “nodes” as follows. “Graph node” refers to an individual entity within a network in the context of graph theory. For example, we may refer to a “graph with 300 nodes” to define the number of nodes within its network. The other use of “node” in this paper is “compute node.” In this work, “compute node” refers to an individual computational unit within the context of a larger HPC environment.
@@ 2.3. Spatiotemporal Data Preprocessing
Figure 1. Simplified workflow of spatiotemporal preprocessing with a $horizon$ of 3. G0 represents the state of the graph at time 0, G1 represents the state of the graph at time 1, and so on to time 4. During preprocessing, a sliding window of size $horizon$ is applied to the data, extracting snapshots. After preprocessing, the data consists of a set of snapshots constructed from the original data.
@@
6:  for each window in $data$ do
@@ 3.2. Results
The PeMS-All-LA dataset presents significant memory challenges for both the DCRNN and PGT-DCRNN implementations. As shown in table 2, DCRNN exhibits high memory consumption, with a peak system memory usage of 371.25 GB and peak GPU memory usage of 24.84 GB. Although still suffering from high memory usage, the PGT-DCRNN model demonstrates substantial lower memory usage (see fig. 2), reducing peak system memory usage to 259.84 GB and peak GPU memory usage to just 1.58 GB. The lower system memory usage in PGT-DCRNN is attributed to standardizing data in place. In contrast, the original DCRNN implementation creates a separate Python object that duplicates the data for standardization, resulting in multiple copies existing in memory during preprocessing. PGT-DCRNN also reduces GPU memory usage and achieves a 15.30x reduction in runtime relative to the original DCRNN model, highlighting its potential as a lightweight foundation for integrating additional scalability techniques.
@@ 3.3. Analysis
Motivated by DCRNN’s high memory usage, we examined the ST-GNN workflow to identify the root causes. Further analysis revealed that the majority of the memory overhead stems from the representation of spatiotemporal graph snapshots produced by SWA. To generate the feature ($x$) and label ($y$) arrays, open-source implementations perform SWA on the source data, extracting an input sequence and label sequence at every valid placement of the window. As shown in fig. 1, preprocessing applies SWA to segment the data into time snapshots (stage 2 in fig. 3). Each of the snapshots contains $horizon-1$ redundant data values, where $horizon$ represents the number of time steps the model is predicting into the future. The data is then further duplicated by creating a $y$ time slice label for each corresponding $x$ time slice (stage 3 in fig. 3). This process and its high degree of data growth are shown graphically for PeMS-All-LA in fig. 3. The resulting data growth can be described analytically as follows: Given a spatiotemporal dataset whose original size is $size=entries\times nodes\times features$, its final size after preprocessing is described by eq. 1, where $entries$ is the total number of dataset items, $horizon$ is the number of time periods the model is predicting into the future, $nodes$ is the number of graph nodes in each graph, and $features$ is the number of data features (e.g., speed, day of week).
[…]
@@ 4.1. Index-Batching
Index-batching modifies ST-GNN preprocessing and training to significantly reduce memory usage. To do so, we take advantage of the unique properties of ST-GNN preprocessing. Given that all data contained in the generated $x$ feature array and the $y$ label array is already present in the original file, we observed that it is more efficient to store only a single copy of the original data (which corresponds to stage 2 in fig. 3) and the graph IDs (see fig. 4) or indices corresponding to the data each $x$-$y$ mapping accesses. Furthermore, since the size of the time slices and the offset between $x$ and $y$ slices are constants defined by the $horizon$, it is unnecessary to store all the indices for each slice or maintain a separate set of indices for the $y$ array. For example, in fig. 4, which uses a $horizon$ of 3, if we store the graph ID that corresponds to the first entry in snapshot 0 (i.e., graph ID 0) in $start$, we can reconstruct the snapshot and its label at runtime as NumPy views by accessing $data[start:start+horizon]$ and $data[start+horizon:start+(2*horizon)]$. This practice removes the high degree of data duplication present in the standard preprocessing workflow by only storing an array of graph IDs roughly equal in size to the number of dataset entries ($\text{entries}-2\times\text{horizon}-1$) and a single copy of the original data. Additionally, by leveraging NumPy views to reference a given snapshot, we avoid copying data during batching. During training, batches are generated on demand by slicing into the standardized data using a set of indices and the $horizon$ value to reconstruct the appropriate snapshots.
@@
Index-batching significantly improves memory efficiency compared with the existing ST-GNN workflow. As shown in eq. 1, the primary factors driving the increase in size during ST-GNN preprocessing are the $horizon$ size and the $2x$ multiplier due to maintaining separate sets of $x$ and $y$ data points. In contrast, index-batching’s space needs are not increased by larger $horizon$ values, and it does not maintain a separate $y$ dataset in addition to the $x$ dataset. Index-batching’s space requirements are represented by eq. 2, where $entries-(2\times horizon-1)$ represents the extra data due to including an array of indices.
[…]
@@ 5. Evaluation
PGT-I has the potential to retain the ease of use of PGT, reduce memory usage in a single-GPU setting, and provide multi-GPU and multinode scalability through specialized distributed training techniques. To evaluate our techniques, we aim to answer three key questions:
@@
Given that existing distributed graph training libraries (e.g., P3 (Gandhi and Iyer, 2021), DGL (Wang et al., 2020), and DIST-DGL (Zheng et al., 2020)) are incompatible with spatiotemporal data, we compare distributed-index-batching with its single-GPU PGT counterpart and a standard DDP workflow, which we refer to as baseline DDP or simply DDP. As described in section 3, PGT represents the state of the art in spatiotemporal training and integrates a range of performance optimizations, making it an ideal single-GPU baseline. Additionally, we compare performance relative to DDP, implemented using PGT for computational efficiency, and Dask.distributed to distribute data across the workers during both preprocessing and training; similar to many large-scale DDP frameworks (Zheng et al., 2020; Elgabli et al., 2020; Shao et al., 2024), this method communicates data on demand across workers. To reduce the cost of communication and ensure a fair comparison, we tested multiple DDP implementations and a variety of Dask configurations (e.g., data transfer block size, threads per worker, spill-to-disk threshold), selecting the best-performing configuration. We found that issuing Dask communication requests for a batch of data rather than individual data items within a batch, which would align with the standard PyTorch sampler’s practice, significantly reduced communication overhead; hence, we incorporated this optimization into our baseline DDP approach. In order to maintain parity with our distributed-index-batching approach and avoid the decreased accuracy associated with local shuffling (Meng et al., 2017), DDP performs global shuffling between epochs.
@@ 5.2. Benefits of GPU-Index-Batching Relative to Index-Batching
As shown in table 4, GPU-index-batching accelerates data preprocessing and model training over index-batching on the PeMS dataset, reducing runtime by 12.87%. Both implementations complete preprocessing in under 30 seconds, with index-batching requiring 26.05 seconds and GPU-index-batching 19.05 seconds. Since preprocessing accounts for only a small fraction of the total training time, we did not prioritize optimizing it for GPU execution. GPU-index-batching’s speedup is due to the elimination of CPU-to-GPU data transfers during training. As a result of storing the dataset in GPU memory, however, GPU-index-batching uses significantly more GPU memory than does index-batching, requiring 18.60 GB relative to index-batching’s 5.50 GB. This is balanced by reducing CPU memory usage by 60.30%. As shown in fig. 6, index-batching causes a spike in CPU memory usage to approximately 46 GB. This initial spike occurs during preprocessing, specifically dataset standardization. In contrast, GPU-index-batching performs this step on the GPU, avoiding the spike and maintaining a lower overall memory footprint.
@@ 5.3.2. Comparison with DDP
Distributed-index-batching significantly outperforms the DDP approach, reducing overall runtime by 2.16x to 11.78x with 4 and 128 GPUs, respectively. This speedup arises from the elimination of inter-node communication and periodic CPU-to-GPU memory transfers, since workers can process data in local memory and utilize a single consolidated CPU-to-GPU memory transfer. As noted in section 5.3.1, because each worker performs preprocessing in local memory, distributed-index-batching’s preprocessing time does not scale with the number of GPUs.
@@ 5.4. Scaling to Larger-than-Memory Datasets
To evaluate the runtime performance of our technique, we utilize the larger PeMS dataset, performing a single epoch of training using 4, 8, 16, 32, 64, and 128 GPUs. We compare with DDP modified to include the changes described above. Notably, while index-batching reduces distributed preprocessing time by an average of 56.86 $\pm$ 30.58 seconds relative to DDP across all configurations, preprocessing constitutes a small portion of the overall runtime in a full training workflow. Therefore, we omit further analysis of preprocessing time.
Figure 9. Batch-shuffling versions of generalized-distributed-index-batching and baseline DDP. Generalized-distributed-index-batching significantly lowers communication cost (as shown by the hashed portion of each bar) by decreasing overall data volume.
Generalized-distributed-index-batching outperforms the baseline epoch time by up to 2.28x. While the baseline’s epoch time improves from 303 seconds using 4 GPUs to 231 seconds, it still suffers from high overhead due to the large volume of communicated data. Generalized-distributed-index-batching also significantly reduces memory usage, using only 53.28 GB of memory with four workers compared with the baseline’s 479.66 GB. Notably, the index-batching implementation’s memory footprint with only 4 GPUs is comparable to the footprint baseline DDP achieved with 32 GPUs (53.30 GB), demonstrating substantial memory savings even in a single-node setting.
@@ 5.5. Broader Applicability
To demonstrate index-batching’s generalizability to sequence-to-sequence models, we perform additional single-GPU experiments with A3T-GCN (Zhu et al., 2020) and the Metr-LA (of California, [n. d.]) dataset. Additionally, we perform a distributed-index-batching scaling study with ST-LLM (Liu et al., 2024) and the PeMS-Bay dataset. A3T-GCN is a spatiotemporal forecasting model that integrates graph convolutions with gated recurrent units and a global attention (Vaswani et al., 2023) mechanism to achieve high accuracy. We integrate index-batching into A3T-GCN’s PGT implementation,  [^6] https://github.com/benedekrozemberczki/pytorch_geometric_temporal/blob/master/examples/recurrent/a3tgcn2_example.py utilizing the provided hyperparameters and training for 30 epochs. ST-LLM utilizes a state-of-the-art approach that encodes spatial-temporal context into token embeddings that are then processed by GPT2 (Radford et al., 2019). ST-LLM is not currently integrated into PGT, and we therefore use its open-source implementation [^7] https://github.com/ChenxiLiu-HNU/ST-LLM. All tests are performed on the Polaris supercomputer, which is described in section 3.1, and we reuse the experimental settings described in section 5. To reduce node-hour cost, we limit the distributed-index-batching tests to 4, 8, 16, and 32 GPUs.
@@ 7. Conclusion
Acknowledgment
This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of Advanced Scientific Computing Research, under grant 0000269227 and contracts DE-AC02-06CH11357, DE-AC02-05CH11231, and DESC0012704. Additionally, this work is supported by NNSA under contract DE-NA0003525.

### v3

@@ 2. Background
This work focuses on the distributed training of ST-GNNs. Accordingly, section 2.1 provides an introduction to spatiotemporal data, section 2.2 introduces ST-GNNs and DCRNN, and section 2.3 describes ST-GNN data preprocessing. To provide clarity, we define the terminology relating to “nodes” as follows. “Graph node” refers to an individual entity within a network in the context of graph theory. For example, we may refer to a “graph with 300 nodes” to define the number of nodes within its network. The other use of “node” in this paper is “compute node.” In this work, “compute node” refers to an individual computational unit within the context of a larger HPC environment.
@@ 2.3. Spatiotemporal Data Preprocessing
Figure 1. Simplified workflow of spatiotemporal preprocessing with a horizonhorizon of 3. G0 represents the state of the graph at time 0, G1 represents the state of the graph at time 1, and so on to time 4. During preprocessing, a sliding window of size horizonhorizon is applied to the data, extracting snapshots. After preprocessing, the data consists of a set of snapshots constructed from the original data.
@@
6:for each window in datadata do
@@ 3.2. Results
The PeMS-All-LA dataset presents significant memory challenges for both the DCRNN and PGT-DCRNN implementations. As shown in table 2, DCRNN exhibits high memory consumption, with a peak system memory usage of 371.25 GB and peak GPU memory usage of 24.84 GB. Although still suffering from high memory usage, the PGT-DCRNN model demonstrates substantial lower memory usage (see fig. 2), reducing peak system memory usage to 259.84 GB and peak GPU memory usage to just 1.58 GB. The higher memory usage in the original DCRNN implementation stems from its custom dataloader, which stores extra copies of the dataset — padded to align with the batch size — in addition to the original data. PGT-DCRNN also reduces GPU memory usage and achieves a 15.30x reduction in runtime relative to the original DCRNN model, highlighting its potential as a lightweight foundation for integrating additional scalability techniques.
@@ 3.3. Analysis
Motivated by DCRNN’s high memory usage, we examined the ST-GNN workflow to identify the root causes. Further analysis revealed that the majority of the memory overhead stems from the representation of spatiotemporal graph snapshots produced by SWA. To generate the feature ($x$) and label ($y$) arrays, open-source implementations perform SWA on the source data, extracting an input sequence and label sequence at every valid placement of the window. As shown in fig. 1, preprocessing applies SWA to segment the data into time snapshots (stage 2 in fig. 3). Each of the snapshots contains horizon1horizon-1 redundant data values, where horizonhorizon represents the number of time steps the model is predicting into the future. The data is then further duplicated by creating a $y$ time slice label for each corresponding $x$ time slice (stage 3 in fig. 3). This process and its high degree of data growth are shown graphically for PeMS-All-LA in fig. 3. The resulting data growth can be described analytically as follows: Given a spatiotemporal dataset whose original size is $size=entries\times nodes\times features$, its final size after preprocessing is described by eq. 1, where entriesentries is the total number of dataset items, horizonhorizon is the number of time periods the model is predicting into the future, nodesnodes is the number of graph nodes in each graph, and featuresfeatures is the number of data features (e.g., speed, day of week).
size $=2\big{[}(\text{entries}-(2\times\text{horizon}-1))$
(1) $\quad\times\text{horizon}\times\text{nodes}\times\text{features}\big{]}$
@@ 4.1. Index-Batching
Index-batching modifies ST-GNN preprocessing and training to significantly reduce memory usage. To do so, we take advantage of the unique properties of ST-GNN preprocessing. Given that all data contained in the generated $x$ feature array and the $y$ label array is already present in the original file, we observed that it is more efficient to store only a single copy of the original data (which corresponds to stage 2 in fig. 3) and the graph IDs (see fig. 4) or indices corresponding to the data each $x$-$y$ mapping accesses. Furthermore, since the size of the time slices and the offset between $x$ and $y$ slices are constants defined by the horizonhorizon, it is unnecessary to store all the indices for each slice or maintain a separate set of indices for the $y$ array. For example, in fig. 4, which uses a horizonhorizon of 3, if we store the graph ID that corresponds to the first entry in snapshot 0 (i.e., graph ID 0) in startstart, we can reconstruct the snapshot and its label at runtime as NumPy views by accessing $data[start:start+horizon]$ and data[start+horizon:start+(2horizon)]data[start+horizon:start+(2*horizon)]. This practice removes the high degree of data duplication present in the standard preprocessing workflow by only storing an array of graph IDs roughly equal in size to the number of dataset entries ($\text{entries}-2\times\text{horizon}-1$) and a single copy of the original data. Additionally, by leveraging NumPy views to reference a given snapshot, we avoid copying data during batching. During training, batches are generated on demand by slicing into the standardized data using a set of indices and the horizonhorizon value to reconstruct the appropriate snapshots.
@@
Index-batching significantly improves memory efficiency compared with the existing ST-GNN workflow. As shown in eq. 1, the primary factors driving the increase in size during ST-GNN preprocessing are the horizonhorizon size and the $2x$ multiplier due to maintaining separate sets of $x$ and $y$ data points. In contrast, index-batching’s space needs are not increased by larger horizonhorizon values, and it does not maintain a separate $y$ dataset in addition to the $x$ dataset. Index-batching’s space requirements are represented by eq. 2, where $entries-(2\times horizon-1)$ represents the extra data due to including an array of indices.
index_batching_size $=\text{entries}\times\text{nodes}\times\text{features}$
@@ 5. Evaluation
PGT-I has the potential to retain the ease of use of PGT, reduce memory usage in a single-GPU setting, and provide multi-GPU and multi-node scalability through specialized distributed training techniques. To evaluate our techniques, we aim to answer three key questions:
@@
Given that existing distributed graph training libraries (e.g., P3 (Gandhi and Iyer, 2021), DGL (Wang et al., 2020), and DIST-DGL (Zheng et al., 2020)) are incompatible with spatiotemporal data, we compare distributed-index-batching with its single-GPU PGT counterpart and a standard DDP workflow, which we refer to as baseline DDP or simply DDP. As described in section 3, PGT represents the state of the art in spatiotemporal training and integrates a range of performance optimizations, making it an ideal single-GPU baseline. Additionally, we compare performance relative to DDP, implemented using PGT for computational efficiency and Dask.distributed to distribute data across the workers during both preprocessing and training; similar to many large-scale DDP frameworks (Zheng et al., 2020; Elgabli et al., 2020; Shao et al., 2024), this method communicates data on demand across workers. To reduce the cost of communication and ensure a fair comparison, we tested multiple DDP implementations and a variety of Dask configurations (e.g., data transfer block size, threads per worker, spill-to-disk threshold), selecting the best-performing configuration. We found that issuing Dask communication requests for a batch of data rather than individual data items within a batch, which would align with the standard PyTorch sampler’s practice, significantly reduced communication overhead; hence, we incorporated this optimization into our baseline DDP approach. In order to maintain parity with our distributed-index-batching approach and avoid the decreased accuracy associated with local shuffling (Meng et al., 2017), DDP performs global shuffling between epochs.
@@ 5.2. Benefits of GPU-Index-Batching Relative to Index-Batching
As shown in table 4, GPU-index-batching accelerates data preprocessing and model training over index-batching on the PeMS dataset, reducing runtime by 12.87%. Both implementations complete preprocessing in under 30 seconds, with index-batching requiring 26.05 seconds and GPU-index-batching 19.05 seconds. Since preprocessing accounts for only a small fraction of the total training time, we did not prioritize optimizing it for GPU execution. GPU-index-batching’s speedup is due to the elimination of CPU-to-GPU data transfers during training. As a result of storing the dataset in GPU memory, however, GPU-index-batching uses significantly more GPU memory than does index-batching, requiring 18.60 GB relative to index-batching’s 5.50 GB. This is balanced by reducing CPU memory usage by 60.30%. As shown in fig. 6, index-batching causes a spike in CPU memory usage to approximately 46 GB. This initial spike occurs during preprocessing, and by performing preprocessing on the GPU, GPU-index-batching reduces the spike and maintains a lower overall memory footprint.
@@ 5.3.2. Comparison with DDP
Distributed-index-batching significantly outperforms the DDP approach, reducing overall runtime by 2.16x and 11.78x with 4 and 128 GPUs, respectively. This speedup arises from the elimination of inter-node communication and periodic CPU-to-GPU memory transfers, since workers can process data in local memory and utilize a single consolidated CPU-to-GPU memory transfer. As noted in section 5.3.1, because each worker performs preprocessing in local memory, distributed-index-batching’s preprocessing time does not scale with the number of GPUs.
@@ 5.4. Scaling to Larger-than-Memory Datasets
To evaluate the runtime performance of our technique, we utilize the larger PeMS dataset, performing a single epoch of training using 4, 8, 16, 32, 64, and 128 GPUs. We compare with DDP modified to include the changes described above. Notably, while index-batching reduces distributed preprocessing time relative to DDP across all configurations, preprocessing constitutes a small portion of the overall runtime in a full training workflow. Therefore, we omit further analysis of preprocessing time.
Figure 9. Batch-shuffling versions of generalized-distributed-index-batching and baseline DDP. Generalized-distributed-index-batching significantly lowers communication cost (as shown by the lighter colored portion of each bar) by decreasing overall data volume.
Generalized-distributed-index-batching outperforms the baseline epoch time by up to 2.28x (see fig. 9). While the baseline’s epoch time improves from 303 seconds using 4 GPUs to 231 seconds, it still suffers from high overhead due to the large volume of communicated data. Generalized-distributed-index-batching also significantly reduces memory usage, using only 53.28 GB of memory with four workers compared with the baseline’s 479.66 GB. Notably, the index-batching implementation’s memory footprint with only 4 GPUs is comparable to the footprint baseline DDP achieved with 32 GPUs (53.30 GB), demonstrating substantial memory savings even in a single-node setting.
@@ 5.5. Broader Applicability
To demonstrate index-batching’s generalizability to sequence-to-sequence models, we perform additional single-GPU experiments with A3T-GCN (Zhu et al., 2020) and the Metr-LA (of California, [n. d.]) dataset. Additionally, we perform a distributed-index-batching scaling study with ST-LLM (Liu et al., 2024) and the PeMS-Bay dataset. A3T-GCN is a spatiotemporal forecasting model that integrates graph convolutions with gated recurrent units and a global attention (Vaswani et al., 2023) mechanism to achieve high accuracy. We integrate index-batching into A3T-GCN’s PGT implementation,  [^6] https://github.com/benedekrozemberczki/pytorch_geometric_temporal/blob/master/examples/recurrent/a3tgcn2_example.py utilizing the provided hyperparameters and training for 30 epochs. ST-LLM utilizes a state-of-the-art approach that encodes spatial-temporal context into token embeddings that are then processed by GPT2 (Radford et al., 2019). ST-LLM is not currently integrated into PGT, and we therefore use its open-source implementation. [^7] https://github.com/ChenxiLiu-HNU/ST-LLM All tests are performed on the Polaris supercomputer, which is described in section 3.1, and we reuse the experimental settings described in section 5. To reduce node-hour cost, we limit the distributed-index-batching tests to 4, 8, 16, and 32 GPUs.
@@ 7. Conclusion
Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of Advanced Scientific Computing Research and the DOE SciDAC program under grant 0000269227 and contracts DE-AC02-06CH11357, DE-AC02-05CH11231, and DESC0012704. Additionally, this work is supported by NNSA under contract DE-NA0003525.